OpenStudy (anonymous):

If you save 1 penny on January 1st, 2 on January 2nd, 3 on January 3rd, and continue this pattern for a year (not a leap year), what will the value of your savings, in dollars, be at the end of one year?

OpenStudy (anonymous):

how many days are in the year without counting the leap year?

OpenStudy (anonymous):

365!

OpenStudy (anonymous):

365 in pennies =

OpenStudy (charlotte123):

A penny everyday - One year - 365 days 1 penny * 365 days = 365 <----

OpenStudy (anonymous):

This is an arithmetic sequence. Try using the following formula.\[1 + 2 + 3 + \cdots + n = \frac{1}{2}n(n+1)\]

OpenStudy (zzr0ck3r):

1+2+3+.... 12 times \[\sum_{i=1}^{12}n\] = 12(12+1)/2

OpenStudy (zzr0ck3r):

@charlotte123 I think you misunderstood the question. you will have 1 at the end of January. 3 at the end of February, 6 at the end of march....

OpenStudy (primeralph):